The mass of a healer and it's contents is 104.334g. You pour 23.5 mL of chloroform into his beaker. What is the mass now? The density of chloroform is 1.49g/mL
1
Expert's answer
2020-01-27T06:22:00-0500

"m=d\times V"

"m(chloroform) = 1.49\frac{g}{mL}\times 23.5 mL = 35.015 g"


"m_{total} = m_{heater} + m_{chloroform} = 104.334 g + 35.015 g = 139.349g"